OpenStudy (anonymous):

Indicate in standard form the equation of the line through the given points. P(0, -4), Q(5, 1)

OpenStudy (anonymous):

do u know the eq of line passing through two points?

OpenStudy (anonymous):

It is as given below: \[\frac{x-x_{1}}{x_{2}-x_{1}} =\frac{y-y_{1}}{y_{2}-y_{1}}\]

OpenStudy (anonymous):

since the line passe sthrough the points P(0, -4), Q(5, 1) therefore \[x_{1}=0, x_{2}=5, y_{1}=-4, y_{2}=1\] plugging these values in the eq we find \[\frac{x-0}{5-0} =\frac{y+4}{1+4} \rightarrow \frac{x}{5} =\frac{y+4}{5}\] i.e. x = y+4 i.e. x-y=4 is the required eq of the line